### v1

# Performances in solving the Bethe-Salpeter equation with the Yambo code

Abstract
In this work, we analyze the performances of two different strategies in solving the structured eigenvalue problem deriving from the Bethe-Salpeter equation (BSE) in condensed matter physics. The first strategy employs direct diagonalization, while the second is based on an iterative solver. The BSE matrix is constructed with the Yambo code, and the two strategies are implemented by interfacing Yambo with the ScaLAPACK and ELPA libraries for direct diagonalization, and with the SLEPc library for the iterative approach. We consider both the hermitian (Tamm-Dancoff approximation) and pseudo-hermitian forms, addressing dense matrices of three different sizes. A description of the implementation is also provided, with details for the pseudo-hermitian case. Timing and memory utilization are analyzed on both CPU and GPU clusters. The CPU simulations are performed on a local cluster in Rome, while the GPU simulations are performed on the Leonardo HPC cluster of CINECA. Our results demonstrate that it is now feasible to handle dense BSE matrices of the order 105.
Keywords:
Eigenvalue problems Bethe-Salpeter Equation GPU
1 Introduction
The Bethe-Salpeter Equation (BSE) is the state-of-the-art first-principles approach to compute neutral excitations in material science [1, 2]. It can be formulated as a non-Hermitian structured eigenvalue problem given by
$\begin{pmatrix}R&C\\ -C^{\dagger}&-A\end{pmatrix}\begin{pmatrix}X_{I}\\ Y_{I}\end{pmatrix}=\omega_{I}\begin{pmatrix}X_{I}\\ Y_{I}\end{pmatrix}$ (1)
where $R$ and $A$ are dense Hermitian matrices, $R=R^{\dagger}$ and $A=A^{\dagger}$, while $C$ is a dense rectangular matrix. Here $R^{\dagger}=(R^{*})^{T}$ denotes conjugate transposition. We refer to the full BSE matrix as $H^{2p}$, and to its structure as pseudo-Hermitian [3, 4, 5, 6]. The BSE structure allows for further simplifications of the eigenvalue problem, allowing more efficient solution methods beyond the standard generalized eigenvalue approach [7]. The same structure also appears in the eigenvalue problems derived within Time-Dependent Density Functional Theory (TD-DFT) and within the Random-Phase Approximation (RPA).
The BSE eigenvalue problem is coded in many condensed-matter (and quantum chemistry) packages, both open-source, such as Yambo [8, 9], BerkeleyGW [2, 10], Abinit [11, 12], Exc [13], Exciting [14, 6, 15], GPAW [16, 17], Simple [18], MolGW [19] and commercial or private, such as VASP [20, 5], TurboMole [21], and Fiesta  [22, 23]. Considering also TD-DFT and RPA, the number of scientific codes implementing the eigenvalue problem of Eq. (1) becomes very large, including applications in quantum chemistry and nuclear physics. Most of them use the standard generalized eigenvalue solvers implemented in eigensolver packages. In condensed matter, most BSE applications are done within the Tamm-Dancoff approximation (TDA) by setting $C=0$, and the BSE eigenvalue problem becomes Hermitian. However, there are physical cases where going beyond the TDA is needed, thus having to deal with the full pseudo-hermitian structure [3, 4, 5, 6].
For the case of optical properties, we have $A=R^{*}$, while $C=C^{T}$. The size of both $R$ and $C$ is $N\times N$. Defining $N_{\mathbf{k}}$ as the number of k-points in the full Brillouin zone, $N_{v}$ as the number of included occupied (valence) states, and $N_{c}$ as the number of unoccupied (conduction) states, $N=N_{\mathbf{k}}\,N_{v}\,N_{c}$ is the total number of optical transitions. Numerically, the most demanding step is the building of the $R$ and $C$ matrices which grows as $\mathcal{O}(N^{2})$, with a rather large prefactor proportional to the (square of the) basis-set size, e.g. squared number of G-vectors, $\mathcal{O}(N_{G}^{2})$, for plane-wave codes. This is why, with the advent of HPC machines and complex structures, a significant effort has been devoted to improve the matrix generation algorithm, taking advantage of MPI, OpenMP and GPU porting schemes such as CUDA, Fortran or OpenACC. Thanks to these developments, matrices with a size of the order $N\propto 10^{5}$ can be easily generated. Even larger matrices can be generated by means of interpolation techniques [24, 25].
The solver step has received significantly less attention, since the computational time needed is usually lower. However, if the solver implementation is not optimized, it can become a barrier, preventing the solution of the eigenvalue problem. The issue is particularly severe when the excitonic wave-functions, e.g. the eigenvectors of the BSE, need to be analyzed. The full diagonalization has a scaling $\mathcal{O}(N^{3})$, and would become the most demanding part of the calculation at a certain size. Iterative schemes show a better scaling, typically $\mathcal{O}(N^{2})$. In both approaches, parallel and GPU ported solvers are needed to handle large matrices. Another issue is the storing of the BSE matrix either on disk (taking advantage of parallel I/O) or directly into RAM memory by distributing them over several nodes/GPU cards. In other eigenvalue problems in condensed matter, the matrix can be easily stored in memory and/or even generated on the fly when using iterative solvers, since (i) the generation process is much faster, and (ii) many matrix elements can be related one to the other. The main example in condensed-matter physics is the eigenvalue problem of DFT, where, as opposed to BSE, most efforts have been devoted to develop efficient and highly parallelized iterative algorithms for the solver step, allowing to reach sizes of the order of $10^{7}$ which are not feasible in BSE. In the BSE instead, explicit storage is needed even if iterative solvers are adopted. Also accounting for some memory redundancy used by the available algorithms, this can easily lead to exceeding the barrier of 1 TB of data to be handled. In the most demanding cases, even tens of TB for a single simulation are required.
In the present work, we analyze the performances of two BSE solvers, one based on direct diagonalization, and the other based on an iterative scheme, and we use them to solve prototype BSE matrices. The two approaches have been implemented in the Yambo code, taking advantage of the interfaces with libraries such as ScaLAPACK [26], ELPA [27, 28], PETSc [29] and SLEPc [30]. We consider both the Hermitian BSE, TDA approximation, and the case in which the full structured pseudo-hermitian matrix is used. For the latter case, we discuss in detail how we take advantage of the pseudo-hermitian structure in both solvers to reduce the computational time. We provide performances on parallel cluster, using a local cluster for CPU only simulations, and the Leonardo machine for GPU ported simulations. An analysis of the memory used on the host is also provided.
2 Description of the algorithm
In this section we provide details about the two computational approaches: full diagonalization and iterative solver. If eigenvectors for more than 5% of the spectrum are required, direct diagonalization routines are recommended. For smaller fractions of the spectrum, iterative solvers are preferable.
2.1 The diagonalization scheme
In this section, we briefly outline the special algorithm for diagonalizing the BSE Hamiltonian given in Eq. (1) when $A=R^{*}$ and $C=C^{T}$, as shown in Ref. [7]. The first step in solving the BSE Hamiltonian $H^{2p}$ with this special structure is to construct a real symmetric matrix $M_{r}$ of the same dimension, which satisfies the relation $Q^{\dagger}H^{2p}Q=-iJM_{r}$, where $J$ is a real skew-symmetric matrix and $Q$ is a unitary matrix:
$Q=\frac{1}{\sqrt{2}}\begin{pmatrix}I&-iI\\ I&iI\end{pmatrix},\quad J=\begin{pmatrix}0&I\\ -I&0\end{pmatrix}.$ (2)
The expression for $M_{r}$ is given by:
$M_{r}=\begin{pmatrix}\text{Re}(R+C)&\text{Im}(R-C)\\ -\text{Im}(R+C)&\text{Re}(R-C)\end{pmatrix}.$ (3)
If $M$ is positive definite, then $M_{r}$ is also positive definite [7], and we can perform a Cholesky decomposition on the real symmetric matrix $M_{r}$, e.g. $M_{r}=LL^{T}$ where $L$ is a real lower triangular matrix. This allows us to cast the original eigenvalue problem into the following form:
$-iWL^{T}\begin{pmatrix}\bar{X}_{I}\\ \bar{Y}_{I}\end{pmatrix}=\omega_{I}L^{T}\begin{pmatrix}\bar{X}_{I}\\ \bar{Y}_{I}\end{pmatrix}$ (4)
where $W=L^{T}JL$ is a real skew-symmetric matrix, and $-iW$ is Hermitian, having the same eigenvalues as the original BSE Hamiltonian. In Eq. (4), the eigenvalues of $W$ differ from the $H^{2p}$ eigenvalues by a factor of $-i$. The eigenvalues of a skew-symmetric matrix come in pairs and are purely imaginary, the eigenvalues of $H^{2p}$ are real and occur in pairs, e.g. $(-\omega_{I},\omega_{I})$. Therefore, the entire BSE problem can be viewed as a real skew-symmetric eigenvalue problem.
In diagonalization, the tridiagonalization step is the computationally expensive part. The algorithm described here enables the tridiagonalization of a real matrix instead of a Hermitian one, significantly reducing both floating-point operations and storage requirements. Once the tridiagonal form of the skew-symmetric matrix is obtained, multiplying the result by $-i$ allows the use of Hermitian tridiagonal solvers to compute the eigenvectors of $W$ [^1] Solving the tridiagonal eigenvalue problem is generally not the primary computational bottleneck, as efficient algorithms are available such as the MRRR method, which scales as $\mathcal{O}(N^{2})$ [31], or the divide-and-conquer approach, which scales as $\mathcal{O}(N^{3})$ but leverages highly optimized General Matrix-Matrix Multiplication (GEMM) operations. After computing the eigenvectors, a back-transformation is applied to retrieve those of the BSE matrix. Moreover, the left eigenvectors for positive eigenvalues, and the left/right eigenvectors corresponding to its negative partner can be determined without explicit computation [7]. More details on the implementation can be found in Ref. [32].
To efficiently interface the algorithm above described with the Yambo code, an MPI redistribution from the Yambo parallel structure to a block-cyclic layout is performed in two main steps. First, the matrix elements are locally grouped along with their corresponding global indices based on their destination rank. Then, a single MPI_AlltoAllv operation is used to transfer the data to their respective destination according to the block-cyclic scheme. Once the matrix is distributed correctly, standard functions from linear algebra libraries can be applied for efficient computations. Here we use ScaLAPACK  [^2] Since ScaLAPACK does not support a real skew-symmetric solver, the tridiagonalization routines for symmetric matrices can be slightly modified to handle skew-symmetric matrices, as both mostly share the same computational operations. and ELPA. For the TDA case, i.e., when the BSE matrix is Hermitian, we use standard Hermitian eigensolvers. ELPA provides two types of Hermitian solver, with type-2 set as the default in Yambo. In ScaLAPACK, the p?heevr solver is used. Both solvers allow one to compute the lowest part of the spectrum. Additionally, ScaLAPACK enables the extraction of eigenvectors within a specified range of eigenvalues or indices.
2.2 The iterative solver
In many cases, a few eigenvalues are sufficient to calculate the optical absorption spectrum with enough accuracy, so an iterative method can be used. Yambo offers two possible approaches: the first is to use a Lanczos iteration with quadrature rules to obtain the optical absorption spectrum directly, without explicitly computing the eigenpairs. The second approach is to use SLEPc to obtain the eigenpairs, as the eigenvectors provide additional information to compute the wave functions. Prior to the release of SLEPc 3.22, the eigenproblem was solved with a standard non-Hermitian two-sided eigensolver, with the possibility to use shift-and-invert to accelerate the convergence of the desired eigenvalues, which are those closest to zero, and lie in the innermost part of the spectrum because of its symmetry. Since the release of SLEPc 3.22, three new methods are available based on [33, 34], which are three variations of a structure-preserving Lanczos iteration with a thick restart.
Essentially, the problem is projected into a smaller structured problem by building a specific Krylov vector basis, which also has a block structure because of the choice of the initial vector. The basis vectors need to be computed accurately and stored in order to obtain the approximate eigenvectors. To prevent the basis from growing too large, a thick restart is used, taking care not to destroy the structure. These methods exploit the structure and properties of the matrix both numerically (symmetry of the matrix blocks and the spectrum) and computationally, working implicitly with the lower blocks of the matrix and vector basis, without storing them explicitly. Additionally, the upper blocks of the matrix can be shell matrices, a type of matrix that does not store its values explicitly, but instead defines how to perform operations with them; in this way, it is possible to reuse Yambo data structures. All three methods are very similar in terms of operations. We focus our performance analysis on the Shao method, which is the current default solver in SLEPc for BSE problems.
The Shao method does not require the use of the computationally expensive shift-and-invert technique because it computes the eigenvalues of $H^{2}$, folding the spectrum so that the internal eigenvalues become external and converge first. A last step is required to obtain the positive eigenvalues and associated eigenvectors of $H$. The negative eigenvalues and their eigenvectors, and the left eigenvectors, can be obtained trivially from them.
MPI parallelization is based on standard row-wise one-dimensional partitioning of matrices and vectors. This is the scheme used by PETSc, which is the library on which SLEPc is built. In the case of GPU simulations, the underlying local computations are offloaded to the NVIDIA GPU using CUDA (or HIP for AMD GPUs). The most expensive step of the algorithm are the matrix-vector products that generate the new vectors of the basis. These vectors need to be generated one by one, as they have to be orthogonalized against the previous ones. Additionally, although the solver is also optimized in the case of a sparse matrix, the problems that arise from applications are typically dense, resulting in an expensive matrix-vector product. The details of the SLEPc solvers for the Bethe–Salpeter eigenproblem are available in [35].
3 Results and discussion
3.1 Details on the performed simulations
Table 1: Characteristics of the HPC clusters
ISM-CNR Cluster Leonardo Booster
CPU model Intel Xeon Gold 5218 Intel Xeon Platinum 8358
CPU Max Freq. 2.30GHz 2.60GHz
Nodes used Up to 3 Up to 32
Cores per node 32 32
Memory per node 86406 [MB] 494000 [MB]
Operating system CentOS Linux Red Hat Enterprise Linux
Release 7.7.1908 Release 8.7 (Ootpa)
GPUs per Node 4
   model NVIDIA A100 GPU HBM2e
   architecture (c.c.) Ampere (8.0)
   memory 64 [GB]
   perf. (S.P.) 75000 [Gflops/node]
CUDA version 11.8
Fortran compiler GFortran 10.2.1 NVFortran 24.9
MPI OpenMPI-4.1.1 OpenMPI-4.1.5
BLAS/LAPACK 3.12 OpenBLAS-0.3.24
ScaLAPACK 2.2.1 2.2.0
ELPA 024.05.001 024.03.001
SLEPc 3.22.0 3.22.1
Yambo 5.3.0 5.3.0
The characteristics of two clusters used for CPU-only and GPU simulations are reported in Table 1. We built in single precision (32-bit float numbers for real and imaginary parts) matrices of Eq. (1) for a single layer of chromium triiodide CrI3. The set of BSE-like matrices considered in the present report are generated by replacing the demanding BSE kernel with the cheaper TD-DFT kernel. We used an energy cut-off of 4 Ry for the eh-exchange $v$, and 10 Ry for the $f_{xc}$ kernel. These parameters determine the pre-factor for the building time of the BSE kernel which we do not analyze here. The size of the BSE matrix is defined by the number of bands used. We considered the three cases reported in Table 2. Here, we focus on the performance of the two solvers described in Sect. 2. For direct diagonalization, ScaLAPACK and ELPA libraries (solver type 2 on CPU, type 1 on GPU) were used with a block size of $64\times 64$ for the block-cyclic distribution. For the iterative solver, the SLEPc library was used, with either explicit (e) or shell (s) matrix formats.
Table 2: Relationship between memory, matrix size and BSE bands. Memory estimated assuming 64-bit single precision floating point complex numbers.
BSE R block Est. Mem. [GB] Type of
bands used size res. case cpl. case Calculation
31-40 10368 0.80 1.60 CPU & GPU
23-40 36288 9.81 19.62 CPU & GPU
23-48 103680 80.09 160.18 GPU only
[…]
  • Strong scaling plots, i.e. time ($t_{i}$) vs number of MPI tasks ($n_{i}$) at fixed matrix size $N_{i}$.
  • Memory scaling plots, i.e. host memory ($M_{i}$) vs number of MPI tasks ($n_{i}$) at fixed matrix size ($N_{i}$).
  • Efficiency plots, i.e. efficiency ($e_{i}(j)$) vs number of MPI tasks ($n_{i}$) at fixed matrix size. For the CPU case $n_{j}=n_{\text{CPU}}=4$, while for the GPU case $n_{j}=n_{\text{GPU}}=2$.
For an efficiency plot, when a reference value is not available, we estimate it using the exponential fit, $t_{i}=t_{1}\,n_{i}^{p}$, where $t_{1}$ represents the time required for a run on a single MPI task, and $p$ is a constant that characterizes how the simulation time scales with the number of processors. For perfect scaling, $p=-1$, i.e., doubling the number of MPI tasks should halve the simulation time. Any value $p>-1$ indicates inefficiencies in parallelization, with a time reduction lower than expected. We will refer to the fitted parameter $p$ as the parallelization factor.
For both CPU and GPU computations, we measured memory usage using the Linux time command, which reports peak memory of each MPI task. For some of the runs we also monitored the peak memory of the run reported by the PETSc internal memory tracking tools. To analyze the memory use, we assume that the allocated host memory is due to arrays/matrices which are either perfectly distributed or fully duplicated. We define the distributed memory as the intercept, and the duplicated memory as the slope of the linear fit of the allocated memory vs number of MPI tasks. Moreover, since in some simulations we observe a sudden memory jump from the serial to the parallel simulations, we also define the parallel–overhead memory as the difference between the fit intercept and the memory allocated in the serial run. We attribute this additional memory increase in parallel runs to the allocation of extra variables not present during serial execution. Memory extracted by the time command and the PETSc tracker agree well, with a parallel overhead detected only by the time command.  [^3] The PETSc data are slightly smaller than the data extracted with the time command in parallel runs. This possibly happens because PETSc tracks the peak memory of the whole run, while time tracks the peak memory of each MPI task, which can be reached at different moments during the simulations.
3.2 CPU scaling. ISM-CNR Cluster
[…]
Table 3: Exponential fit of time vs number of CPUs, $t_{i}=t_{1}\,n_{i}^{p}$. Where $p=-1$ is the ideal parallelization case. (e) - explicit matrix type, (s) - shell matrix type.
$p$ (Fitted Exp.) $p$ (Fitted Exp.)
Solver N=10368 N=36288 Solver N=10368 N=36288
 ScaLAPACK res. -0.74 -0.74  SLEPc res. (e) -0.63 -0.65
 ScaLAPACK cpl. -0.79 -0.77  SLEPc cpl. (e) -0.67 -0.66
ELPA res. -0.87 -0.82  SLEPc res. (s) -0.25 -0.66
ELPA cpl. -0.82 -0.94  SLEPc cpl. (s) -0.36 -0.60
Table 4: Distributed (Distr.), duplicated (Dupl.), and parallel overhead (Par.) memory for the CPU case. Memory tracked with time command (PETSc tracker in parenthesis). (e) - explicit matrix type, (s) - shell matrix type. All units in [GB].
N=10368 N=36288
Solver Distr. Dupl.   Par. Distr. Dupl.   Par.
ScaLAPACK res. 4.96 0.37 -0.19 55.82 0.47 -
ScaLAPACK cpl. 13.54 0.45 5.96 205.98 0.34 -
ELPA res. 4.97 0.41 -0.14 68.83 0.37 -
ELPA cpl. 16.15 0.42 -0.47 208.09 0.32 -
SLEPc res. (e) 3.68 (0.58) 0.42 (0.41) 4.15   -   (9.95)   - (0.44) -
SLEPc cpl. (s) 4.68 (0.80) 0.53 (0.51) 4.82   - (17.56)   - (0.60) -
SLEPc res. (s) 0.15 (0.15) 0.39 (0.40) -0.39   -   (5.73)   - (0.42) -
SLEPc cpl. (s) 0.21 (0.22) 0.46 (0.47) -1.00   - (11.31)   - (0.53) -
Fig. 1 shows that the ELPA library outperforms ScaLAPACK library. This is also reflected in the p values reported in Table 3, which are closer to -1 for ELPA. In fact, among all the studied solvers, ELPA demonstrates one of the best efficiencies and scaling behaviors. As the matrix size increases, ELPA’s efficiency will plateau depending on the size of matrix and type of matrix, with larger matrices benefiting more from ELPA’s optimization, whereas for the ScaLAPACK the plateau always happens around the value of efficiency in the range 0.4 - 0.5.
Figure 1: Execution time, memory usage and efficiency plots by different solvers and different matrices. CPU case.
SLEPc generally runs about an order of magnitude faster than its full diagonalization counterparts, here we ask to extract 100 eigenvalues. The efficiency of the SLEPc solver is in general worse than the diagonalization case. The shell matrix algorithm, in particular, performs very poorly, with a total time which starts to grow after a critical number of MPI tasks is reached. This is however, in part, expected. Since in the shell approach the algorithm uses a matrix distribution which is optimized for building of the kernel. The explicit approach is instead more efficient, although not as efficient as full diagonalization.
As shown in Fig. 1, efficiency drops from 1 to about 0.4 when using up to 16 cores. Beyond that point, efficiency plateaus for the explicit case (while continuing to decline for the shell case). Additionally, for SLEPc, neither the type nor the size of the matrix significantly affects efficiency, which consistently plateaus around 0.4.
Duplicated memory remains consistently around 0.4–0.5 GB per CPU across all cases, likely due to internal Yambo operations and is independent of the solver. Its impact decreases with matrix size. In contrast, distributed memory depends on both solver and matrix size. Direct diagonalization shows high memory usage – 5–6 times larger than the calculated value for resonant case (Table 2); and coupling case uses about 3-4 times more memory than the resonant case. SLEPc requires much less memory, enabling single MPI runs, unlike full diagonalization. In the explicit case, memory usage increases notably when transitioning from serial to parallel runs which we attribute to parallel overhead (see Table 4). This effect is likely due to additional variable allocations. Reducing this overhead could enhance SLEPc’s efficiency.
3.3 GPU scaling. Leonardo Booster
Figure 2: Execution time, peak memory, and efficiency plots by different solvers on matrices of 3 different sizes. GPU case. For SLEPc, coupling matrix, data extrapolated assuming $\mathrm{time}=\frac{4973}{1.7^{\log_{2}(\mathrm{nr\_gpu})}}$. Only host memory is reported here.
Table 5: Exponential fit of time vs number of GPUs, $t_{i}=t_{1}\,n_{i}^{p}$. $p=-1$ is the ideal parallelization case.
$p$ (Fitted Exponential)
Solver  N=10368  N=36288  N=103680
 ELPA res. -0.17 -0.62 -0.67
 ELPA cpl. -0.70 -0.93 -0.91
 SLEPc res. -0.34 -0.65 -0.87
 SLEPc cpl. -0.37 -0.63 -0.40
Table 6: Distributed (Distr.) and duplicated (Dupl.), and parallel overhead (Par.) memory for the GPU case. All units in [GB].
N=10368 N=36288 N=103680
  Solver   Distr.   Dupl.    Par.   Distr.   Dupl.   Par.   Distr.   Dupl.
ELPA res 5.02 2.08 -0.16 64 2.00 -0.17 526 2.00
ELPA cpl 16.51 2.03 -0.38 208 1.98 - 1720 1.85
  SLEPc res 4.13 2.11 4.36 72 2.45 63.92 861 1.95
SLEPc cpl 6.32 2.15 5.68 102 3.45 71.89 1172 2.26
[…]
Medium-sized matrices are handled more efficiently. Their parallelization coefficient is higher (closer to the values observed in the CPU case, see Tables 3 and 5). For these matrices, a performance plateau is reached around 32 GPUs for the resonant case and around 64 GPUs for the coupling case. Up to the plateau, the scaling is very good: doubling the number of GPUs roughly halves the execution time, though the plateau is reached relatively quickly.
For the big matrices the scaling is even better. This is evident from the efficiency plots in Fig. 2, and from the parallelization factors in Table 5. We do not observe any plateau up to the maximum number of GPU used, e.g. 64 (128) for the resonant (coupling) case. Overall we can conclude that, the larger the matrices, the better the scaling and the effciency.
Similar to the CPU case, SLEPc executes an order or two of magnitude faster than ELPA. For the SLEPc solver oscillations in the execution time have been observed. No consistent pattern was found, but data suggest that numbers of GPU equal to powers of $2$ perform better.
Memory usage analysis shows that duplicated memory remains consistently around 2 GB per GPU, which is 4 times higher than for the CPU case. The distributed memory is very close to the one obtained in the CPU case, see Table 4. On Leonardo Booster, per each GPU there is 64GB of memory, and for the cases of medium and big matrices, this imposes a minimal amount of GPUs to perform the simulations. In comparison with the CPU case, on GPUs the SLEPc solver shows similar or even higher memory consumption than the ELPA solver. This is due to the large parallel overhead that we are presently investigating.
For the GPU implementation, the parallel overhead is of a similar order of magnitude to that observed in the CPU case. Specifically, for SLEPc, the parallel overhead is large and positive, indicating the allocation of additional variables for parallel runs. Conversely, for ELPA, the parallel overhead is negative, suggesting more efficient parallelization schemes.
4 Conclusions
We have shown that solvers based on ELPA, ScaLAPACK and SLEPc libraries can efficiently handle very large BSE matrices, both in the resonant and in the coupling case. This is possible since the interface with the libraries, as implemented in the Yambo code, distributes the memory load over the MPI tasks. Large matrices can then be handled, provided that a sufficient number of MPI tasks is used. Moreover, we have shown a good scaling both on CPUs and GPUs (for ELPA and SLEPc). For exact diagonalization, the ELPA library outperforms the ScaLAPACK library on CPUs. For the iterative solver based on SLEPc, it is more efficient to redistribute the matrix in memory from the Yambo parallelization scheme to the SLEPc parallelization scheme (explicit solver) at the price of some memory duplication. Re-distribution in memory is also needed for ELPA and ScaLAPACK.
In conclusion, solvers based on external libraries can be reliably used for the Bethe-Salpeter equation (BSE). The situation is somewhat different from other cases, such as the DFT eigenvalue problem, where the solver is one of the most demanding steps. Accordingly, for the DFT case, it might be better to have DFT-specific and more efficient iterative solvers directly inside the code. For the BSE case, the building of the matrix remains, in many cases, the most demanding step on which code developers can focus. Therefore, opting for libraries-based solvers is a valuable solution, because it allows us to overcome the “solver barrier”, which appears for very large matrices, especially when extraction of the eigenvectors is needed. The biggest advantage for BSE code developers in using libraries is that they outsource the implementation, MPI distribution, and GPU (or other architecture) porting. This approach is well justified by the performance results presented in this work.
4.0.1 Acknowledgements
Innovation Study ISOLV-BSE has received funding through the Inno4scale project, which is funded by the European High-Performance Computing Joint Undertaking (JU) under Grant Agreement No 101118139. The JU receives support from the European Union’s Horizon Europe Programme. D.S. acknowledges funding from the MaX “MAterials design at the eXascale” project, co-funded by the European High Performance Computing joint Undertaking (JU) and participating countries (Grant Agreement No. 101093374). D.S. also acknowledges funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement 101118915 (project TIMES). D.S. and P.M. acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support, through the project IsCc6_ISOL-RUN. B.M., F.A., E.R., and J.E.R. were partly supported by grant PID2022-139568NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. B.M was also supported by Universitat Politècnica de València in its PAID-01-23 programme. NM and LW acknowledge the use of HPC facilities of the University of Luxembourg.
Disclosure of Interests. The authors have no competing interests to declare that are relevant to the content of this article.